OpenStudy (anonymous):

can someone explain this problem and show workplease. 6(3-z) over 5 = -z

OpenStudy (turingtest):

do you mean [6(3-z)]/5=-z ?

OpenStudy (anonymous):

no......6(3 - z)/5=-z......this is exactly the way it is written in by book.

OpenStudy (turingtest):

that's the same I guess, so multiply both sides by 5: 6(3-z)=-5z distribute on the left: 18-6z=-5z add 6z to both sides and combine like terms: 18=6z-5z=z therefor z=18

OpenStudy (anonymous):

Thanks!